OpenStudy (anonymous):

Sin(pi/15)Cos(4pi/15)+Cos(pi/15)Sin(4pi/15)

OpenStudy (anonymous):

this is the addition angle formula for sine since \[\frac{1}{15}+\frac{4}{15}=\frac{5}{15}=\frac{1}{3}\] your answer is \[\sin(\frac{\pi}{3})\]

OpenStudy (anonymous):

Why can you just drop the pi?

OpenStudy (anonymous):

i didn't drop the \(\pi\) it is right here in \[\sin\left(\frac{\pi}{3}\right)\]

OpenStudy (anonymous):

oh, okay! Thank you!